OpenStudy (anonymous):

2sin^4(z) + 7cos(z) = 4, between 0 and 360 inclusive. i could turn sin^4(z) into 1-cos^4(z), and then turn that into (1+cos^2(z))(1-cos^2(z)), but then what??

hartnn (hartnn):

how is sin^4(z) = 1-cos^4(z) ??

OpenStudy (anonymous):

isn't it??

hartnn (hartnn):

nopes

OpenStudy (anonymous):

then what can i do here??

hartnn (hartnn):

maybe (1-cos^2 x)^2 = (1- 2 cos^2 x + cos^4 x) (i have used sin^2 + cos^2= 1, so sin^2 = 1-cos^2)

hartnn (hartnn):

2-4y^2+2y^4+7y=4.......by putting y=cos x not easy to solve, can u use calculator?

hartnn (hartnn):

am i making any sense?

hartnn (hartnn):

@estudier any ideas?

OpenStudy (anonymous):

Just looking....

OpenStudy (anonymous):

Where is the question from?

OpenStudy (anonymous):

cos z= (1+ sin 2z)/2 and let's say sin z= y ..you are done

OpenStudy (anonymous):

i mean cos^2 z

hartnn (hartnn):

trignometry only, i guess

OpenStudy (anonymous):

sin^2 z=y

hartnn (hartnn):

how would that help @jasonxx ? what about sin 2x then?

OpenStudy (anonymous):

it'll be a quadratic equation and it is done

OpenStudy (anonymous):

something like 2y^2.....

hartnn (hartnn):

quadratic?? u have sin2x not sin^2 x

OpenStudy (anonymous):

wait a sec guys ..we all are confused the question is

OpenStudy (anonymous):

\[2 \sin^4 z+ 7 \cos z=4\]

OpenStudy (anonymous):

\[\cos z= \sqrt {1- \sin^2 z}\] let's say \[\sin^2 z= y\] now i think it can be done

hartnn (hartnn):

u just increased the complexity by bringing square root sign....

OpenStudy (anonymous):

\[2y^2+7(\sqrt{1-y})=4\] \[2y^2-4=7 \sqrt{1-y}\] square both sides and then let's say y^2=z

OpenStudy (anonymous):

yea ..there could be alternative

hartnn (hartnn):

so u will get the same equation with y^4 which i hace posted

hartnn (hartnn):

2-4y^2+2y^4+7y=4

OpenStudy (anonymous):

\[4y^4-8y^2+16=49(1-y)\]

hartnn (hartnn):

even that is not any simpler to solve

OpenStudy (anonymous):

@tanvidais13 stuck or done lol, because you seem to be frozen ;)

OpenStudy (anonymous):

sorry, guests over, had to "entertain" :P

OpenStudy (anonymous):

are you sure you wrote the question correctly ?

OpenStudy (anonymous):

no it is done wait a sec

OpenStudy (anonymous):

\[\sin^4 z= (\sin^2z)^2= (1-\cos^2z)^2\] now \[1-2\cos^2z+\cos^4z=(1-\cos^2z)^2\]

OpenStudy (anonymous):

what jasonxx wrote in the beginning is right.

OpenStudy (anonymous):

@tanvidais13 hope i entertained lol

hartnn (hartnn):

didn't entertain, didn't solve.... and the last thing u did,was the 1st thing i did

OpenStudy (anonymous):

you are writing the same thing again

OpenStudy (anonymous):

wait i STILL didn't get how you got 1−2cos^2(z)+cos^4(z)=(1−cos^2(z))2 what is that???

OpenStudy (anonymous):

@hartnn i know you might be fillipn out..relax mate, try to see things from everyone's visual angel

OpenStudy (anonymous):

(a-b)^2 = a^2 -2ab + b^2

OpenStudy (anonymous):

oh yeahh, so that was just the expansion.

hartnn (hartnn):

\(sin^2 x+cos^2 x =1\)and that expansion

OpenStudy (anonymous):

@tanvidais13 i wrote sin^2 z = 1- cos^ z and that was an expansion

OpenStudy (anonymous):

\[(1- \cos^2 z)= \sin^2 z\] and \[(1-\cos^2z)^2= expansion (a-b)^2\] ;)

OpenStudy (anonymous):

but then we only get cosz in powers of 1,2, and 4, not three. can we still solve it like that??

OpenStudy (anonymous):

yes you can...with a basic rearrangement

OpenStudy (anonymous):

ENTERTAIN ME :P

OpenStudy (anonymous):

are you sure ? lol

OpenStudy (anonymous):

\[\cos^4z-2\cos^2z+1+7cosz=4\] now can you play with this equation , see something is common among them, which can be shotned

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

\[2 \cos^4z-4\cos^2z+2+7\cos z =4\]

mathslover (mathslover):

@jasonxx I request you to please don't interrupt when some one is helping a user. No worries, if you were ONLY giving suggestion but if you are spoiling the work of the previous user then it will not be nice.

mathslover (mathslover):

Do refer : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@mathslover i do respect your words but i can't stop myself from some arrogant person and his haughty comment i am sure you know, whom am i talking about

OpenStudy (anonymous):

if we say cos(z) = y, then we will have \[2y^4 - 4y^2 + 7y - 2 = 0\]

OpenStudy (anonymous):

@tanvidais13 \[2y^4-4y^2=2-7y\] now? can you do the rest

mathslover (mathslover):

@jasonxx please stay friendly dude. It would be better if you have not used "arrogant" like words, this makes the whole difference. It's nice to see that you respect my words but I will say that even I am spoiling this post now, so I am leaving , you may continue your help but remember the following things from now onwards: i) Never give direct answer to any asker. ii) Don't give direct solution to the asker, this not only ends the motive of openstudy but also not helps asker. iii) Let the asker do the work, you must suggest him/her regarding your idea or steps. like : 5x-7=3 , just say "Try to add 7 both sides, what do you get?" iv) "Answer snipping" , this is what you have likely done here, if a user is helping the asker then please don't interrupt by giving direct solution or any way, but if you have good ideas then just post shortly. Thanks :)

hartnn (hartnn):

see my 4th comment, 35 minutes ago, we are back there, we wasted 35 minutes..

OpenStudy (anonymous):

@mathslover thank you but you should not be biased..well for the sake of us i am putting this to an end, i haven't copied i was friendly and i was working the way i believe is dood and is in accordance with this code of conduct but if one is willing to have my welcome in some other way then i am sorry, either i'll help or i shall feel compelled to assault someone verbally ...my serious apology for being rude thank you

mathslover (mathslover):

Thanks and sorry if i was rude, but now someone help @tanvidais13 :)

OpenStudy (anonymous):

@Mikael have a look, i believe you can take us outta this

OpenStudy (anonymous):

@tanvidais13 if we take value of Z=0 degrees ... this equation fails !! are you sure this question is correct?

OpenStudy (anonymous):

2*0+7*1 can't be equal to 4 ?? isn't it?

OpenStudy (anonymous):

@mathslover what say bro? :) any idea !!

mathslover (mathslover):

not yet..

mathslover (mathslover):

@sauravshakya and @UnkleRhaukus

OpenStudy (anonymous):

doesn't it look like equation that is not consistent ..it fails whilst substituting the value of z

OpenStudy (anonymous):

but you had to find z .. you dont plug any value you want

OpenStudy (anonymous):

good point, but how one is supposed to get value of somethin from (not sure) an incorrect equation

OpenStudy (anonymous):

who said that it is incorrect

OpenStudy (anonymous):

intuition ...let me solve this again :)

OpenStudy (anonymous):

Cant this be factored: